At the beginning of the year, Grillo Industries bought three used machines from Freeman Incorporated. The machines immediately were overhauled, were installed, and started operating. Because the machines were different, each was recorded separately in the accounts.

Machine A Machine B Machine C
Cost of the asset $ 9,800 $ 39,000 $ 22,800
Installation costs 950 2,900 2,000
Renovation costs prior to use 750 2,500 3,000
Repairs after production began 500 900 1,500

By the end of the first year, each machine had been operating 8,000 hours.

Required:

  1. Compute the cost of each machine.
  2. Prepare the journal entry to record depreciation expense at the end of year 1, assuming the following:
Estimates
Machine Life Residual Value Depreciation Method
A 5 years $ 1,800 Straight-line
B 20,000 hours 1,400 Units-of-production
C 10 years 2,200 Double-declining-balance

Complete this question by entering your answers in the tabs below.

  • Required 1
  • Required 2

Prepare the journal entry to record depreciation expense at the end of year 1, assuming the following: (If no entry is required for a transaction/event, select "No Journal Entry Required" in the first account field. Do not round intermediate calculations.)

No Transaction General Journal Debit Credit
1 1 Depreciation Expense
Accumulated Depreciation - Machine A 3,980
Accumulated Depreciation - Machine B
Accumulated Depreciation - Machine C
  • Required 1
Answer #1

Requirement 1

Total Cost of Machine
Machine A $      11,500.00
Machine B $      44,400.00
Machine C $      27,800.00

Working

Amount paid for assets Installation cost Renovation cost Total Cost of Machine
Machine A $         9,800.00 $            950.00 $               750.00 $    11,500.00
Machine B $       39,000.00 $        2,900.00 $           2,500.00 $    44,400.00
Machine C $       22,800.00 $        2,000.00 $           3,000.00 $    27,800.00

Requirement 2

Transaction General Journal Debit Credit
1.. Depreciation expense $    24,700.00
           Accumulated depreciation-Machine A $        1,940.00
           Accumulated depreciation-Machine B $      17,200.00
           Accumulated depreciation-Machine C $        5,560.00
(To record depreciation expense)

Working

Straight line Method
A Cost $ 11,500
B Residual Value $ 1,800
C=A - B Depreciable base $ 9,700
D Life [in years left ]                                  5
E=C/D Annual SLM depreciation $               1,940.00

.

Units of Production method
A Cost $ 44,400
B Residual Value $ 1,400
C=A - B Depreciable base $ 43,000
D Usage in units(in Hours) 20000
E Depreciation per hour $                        2.15

.

Depreciation schedule-Units of Activity
Year Book Value Usage Depreciation expense Ending Book Value Accumulated Depreciation
1 $                44,400 8000 $                   17,200 $                    27,200 $               17,200

.

Double declining Method
A Cost $ 27,800
B Residual Value $ 1,500
C=A - B Depreciable base $ 26,300
D Life [in years] 10
E=C/D Annual SLM depreciation $ 2,630
F=E/C SLM Rate 10.00%
G=F x 2 DDB Rate 20.00%

.

Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
1 $ 27,800 20.00% $ 5,560 $ 5,560 $ 22,240